OpenStudy (anonymous):

8th grade math... Find two consecutive odd integer whos sum is 28 please explain how u got ur answer

OpenStudy (saifoo.khan):

x + (x+2) = 28 Solve for x.

OpenStudy (anonymous):

u have to set up an equation x+ x+2 = 28 it becomes 2x+2 = 28. Isolate x and u get 13 as ur fist consecutive integer and 15 as ur second consecurtive integer

OpenStudy (anonymous):

15 is the 2nd odd integer since the second number is x+2

OpenStudy (anonymous):

give me a good answer sicne i explained it

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

no problem